# Federated Learning for Smart Grid: A Survey on Applications and Potential Vulnerabilities

### v1

@@ 1. Introduction
The Smart Grid (SG), which integrates the physical infrastructure of the electrical grid with modern information and communication technologies (ICTs) (Vasisht et al., 2022), is a prime example of a cyber-physical system (CPS). Unlike conventional power grids, SGs can detect, respond to, and proactively manage changes in electricity demand and supply, while also addressing various operational challenges (Obeagu and Owunna, 2022).
The SG architecture incorporates several key components that work together to enhance the overall functionality of the electrical system (Mittal et al., 2024), including advanced metering infrastructure (AMI) (Molokomme et al., 2019), communication networks (Hu et al., 2023), and sophisticated information technology solutions (Reza et al., 2014), among others. By leveraging real-time data and advanced analytics, SGs aim to improve the efficiency of power generation, transmission and distribution, and consumption (Chopda et al., 2021). Although SGs enable a more flexible and resilient power system that can better accommodate the growing demand for electricity, the increasing reliance on ICTs has made SG systems more vulnerable to security and privacy attacks. Federated Learning (FL) has emerged as a promising approach to addressing privacy and security concerns in SG systems. This innovative distributed machine learning paradigm enables collaborative model training across multiple participants without the need to share raw data, making it particularly suitable for SG applications, where generators are distributed and data privacy for each generator is crucial. We categorize the application of FL in SGs into three aspects: generation, transmission and distribution, and consumption. From the perspective of generation (Tuballa and Abundo, 2016) in SGs, FL is primarily being used to enhance renewable energy integration and forecasting (Camacho et al., 2011). FL enables distributed SG generators to collaboratively train models for predicting resources such as solar and wind power output without sharing sensitive operational data. In the transmission (Bayoumi, 2015) and distribution aspects (Hamidi et al., 2010) of SGs, FL is proving valuable for line fault detection and classification, optimal power flow (Usman and Capitanescu, 2022), voltage control (Xie et al., 2021), and more. By utilizing a FL-based approach, grid operators can quickly identify and locate faults in transmission lines, while collaboratively training models for power distribution without compromising the privacy of individual generators. In the consumption process of SGs (Kaur et al., 2020), FL is extensively used for load forecasting (Husnoo et al., 2023) and energy theft detection (Li et al., 2024a). Smart meters can collaboratively train models to predict household energy consumption patterns without sharing individual usage data.
@@
Grataloup et al. (Grataloup et al., 2024) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zheng et al. (Zheng et al., 2024) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zhang et al. (Zhang et al., 2024) 2024 $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Pandya et al. (Pandya et al., 2023) 2023 $\surd$ $\surd$ $\surd$ $\surd$
Cheng et al. (Cheng et al., 2022) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Mirzaee et al. (Mirzaee et al., 2022) 2022 $\surd$ $\surd$ $\surd$
Ding et al. (Ding et al., 2022) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Cui et al. (Cui et al., 2020) 2020 $\surd$ $\surd$ $\surd$ $\surd$
@@ 1.1. Related Surveys
Grataloup et al. (Grataloup et al., 2024) focus on the integration of FL to address challenges in renewable energy applications, such as energy management, system reliability, energy consumption prediction, and load forecasting. However, this survey only highlights FL use cases in SGs as privacy-preserving collaborative learning approaches and overlooks the potential vulnerabilities of SGs when using deep learning and federated learning technologies. Zheng et al. (Zheng et al., 2024) provide a detailed discussion of components in FL systems, including model aggregation, data partitioning, and local model training. However, their survey paper lacks a discussion on model-related attacks within FL-based SG systems. Zhang et al. (Zhang et al., 2024) offer a comprehensive survey of applications for each stage of SG systems and discuss potential privacy and security violations. However, their focus is primarily on centralized machine learning algorithms. Our work bridges the gaps between SG, FL, and privacy and security vulnerabilities. Pandya et al. (Pandya et al., 2023) demonstrate a broader scope, covering smart cities with a FL framework to preserve privacy. This approach enables effective utilization of large-scale data in areas such as electric supply chain monitoring, air quality forecasting, water consumption prediction, traffic management, and pandemic diagnosis. However, this work lacks details on SG applications and security issues within FL frameworks. Cheng et al. (Cheng et al., 2022) conduct a survey of the FL paradigm, comparing the types, network structures, and global model aggregation methods in SG. However, the brief survey lacks a discussion on robustness. Mirzaee et al. (Mirzaee et al., 2022) and Ding et al. (Ding et al., 2022) discuss conventional security threats to SGs as well as emerging issues related to the application of machine learning techniques. They cover various types of cyberattacks, privacy concerns related to smart meter data, and potential countermeasures to address these challenges. However, they only cover a limited number of applications in SGs and primarily focus on centralized cases. Cui et al. (Cui et al., 2020) mainly focus on false data attacks and privacy issues in SGs. Compared to previous literature, our paper provides a comprehensive discussion on FL-based SGs and their potential vulnerabilities. We hope this can complement existing research and inspire further studies in this direction.
@@ 2.1. Smart Grid
The electric power grid has evolved by leaps and bounds since its inception (Masera et al., 2018). A major development in this domain was the integration of ICT. The integration of ICT with the traditional power grid facilitated distributed power generation via better resource management and automated control (Simões et al., 2012). ICT integration also marked the evolution of the conventional grid into the SG and solved several limitations including unidirectional power flow and centralized bulk generation of electricity. Additionally, it reduced the power and energy domain’s dependence on fossil fuels like coal and natural gas, leading the way for sustainability and decarbonization via the development of microgrids (Jahromi et al., 2023; Yadav et al., 2020). The SG power flow framework consists of the following stages:
@@
    Generation: This stage refers to the production of AC/DC power. In the conventional power grid, power generation was centralized in nature. However, with ICT integration for real-time monitoring, power generation can also be achieved in a regulated manner at residential buildings (for example using solar photovoltaic panels) and used for local consumption/fed back to the grid (Kahrobaee et al., 2012).
@@
    Transmission/Distribution: This stage refers to the transfer of the power produced at the generation stage to the appropriate end nodes for utilization. This involves the use of transformers, transmission lines, substation equipment and other line devices like fault detector relays, breakers to ensure a safe and reliable transfer at appropriate bus voltage levels. The fault detector relays and breakers are often digital in nature and controlled by the central monitoring station via ICT devices (Gaspar et al., 2023).
@@
    Consumption: This is final stage of the SG power flow process and involves the utilization of the power as received from the transmission and distribution networks for powering electrical devices. The consumption level can involve entities like government agencies, private industries, private individuals, etc., each of which can have a different consumption footprint. In the modern SG, integration of ICT devices means that some consumer-level devices like electric vehicles can also be used to transfer surplus/unused power back to the grid (coordinated via V2G communications (Madawala and Thrimawithana, 2011)).
@@
While ICT integration in the power grid enabled easier and more robust resource management from a single centralized Supervisory Control and Data Acquisition (SCADA)-based control system (Gaushell and Block, 1993), the increasing number of IoT devices and local generation facilities like solar panels means that it can be extremely computationally resource intensive to monitor them. Inadequate monitoring facilities can mean that intrusive events like cyberattacks and line faults may not be detected and mitigated on time.
@@
Slowed mitigation of cyberattacks, faults and other control emergencies can have a devastating impact on the stability of the SG (Jahromi et al., 2023). Moreover, the SG being a cyber-physical system implies that cyber layer vulnerabilities can be exploited to cause severe physical ramifications. For example, cyberattacks on the SG can lead to uncontrolled voltage increase at one or more buses leading to equipment damage and consumer-level device burnout (Rath et al., 2022). To alleviate these issues, several researchers have proposed, implemented and evaluated the use of Artificial Intelligence (AI)-based tools and techniques for applications like real-time control (Bose, 2017), cybersecurity (Ruan et al., 2023), load forecasting (Zhou and Jin, 2017), and fault detection (Sun et al., 2023). These AI tools are typically deployed at the centralized controller level for monitoring, removing the need for human system operators.
@@
As the AI modules employed for applications like cybersecurity/control monitoring and load forecasting in the SG environment, are often placed at a central server, they require the exchange of data between local Distributed Energy Resources (DERs) and the central node via communication networks. However, this presents privacy risks (Sun et al., 2021) as the data being transferred can be utilized to reveal energy consumption patterns of private individuals (Liu et al., 2018; Zhao et al., 2020; Wang et al., 2020b). This is a major risk as the consumption patterns can be cross-analyzed to determine when the individuals are at home/outside. Further, their home appliances can also be revealed motivating malicious entities to schedule burglary attempts (Liu et al., 2018).
@@
Another risk that the communication-based training data exchange between local DERs and the centralized server housing AI modules is the non-willingness within third-party DER owners to share proprietary data. This non-willingness typically stems from the concern that the proprietary data may be accessed by malicious entities who can then know the system details of the DER. This can be used to manipulate the system via stealth cyberattack vectors (Leng et al., 2021). Non-willingness can also arise from a lack of trust between the DER owners and the utility company that manages the SG and the control center.
@@
Both the above-mentioned factors can lead to data constraints that may severely hinder the AI training process leading to flawed and untrustworthy decision-making, raising concerns on the reliability of AI. To solve the problem, An AI training strategy called Federated Learning can be used. In this strategy proprietary training data does not need to be transferred over a communication channel leading to alleviation of privacy concerns (McMahan et al., 2017a). More details about this training strategy are available in the following subsection.
@@ 2.2. Federated Learning
To solve the above optimization problem, the system runs for $T$ rounds of FL training. Initially, the server stores the global model $\mathbf{w}^{0}$. In round $t$, the server randomly selects a subset $\mathcal{Q}^{t}$ of $q$ devices and sends them the latest global model $\mathbf{w}^{t}$. These selected devices update the global model $\mathbf{w}^{t}$ using their local data. For instance, in the widely-used Federated Averaging (FedAvg) (McMahan et al., 2017b) algorithm, each selected device performs $\kappa$ iterations of stochastic gradient descent (SGD) to update the local model as follows:
@@ 3.1. Generation Stage
FL enables collaborative model training across different energy producers and consumers, enhancing the accuracy of energy generation forecasts and enabling dynamic resource allocation without the need to centralize sensitive operational data (Lee, 2023; Gupta et al., 2023). In the context of energy management (Abdellatif et al., 2024), FL supports the development of intelligent systems that can predict and optimize power generation patterns while keeping sensitive user data local on smart devices. This approach improves grid efficiency and facilitates more effective demand response programs.
@@
Energy Forecasting In the work by (Lee, 2023), an advanced algorithm for FL is proposed to tackle the challenges of non-IID data in distributed solar energy grids. This framework is specifically designed to handle the heterogeneous nature of the data across different solar energy producers, enhancing the model’s ability to converge efficiently in decentralized settings. By addressing the variations in data distribution, the framework improves the accuracy and reliability of predictions, leading to better performance in solar energy generation and more effective resource allocation.
Energy Optimization The FedGrid framework introduced by (Gupta et al., 2023) offers a secure FL solution for optimizing energy management within smart grids. This framework emphasizes the importance of protecting privacy and enhancing data security during the optimization process. By enabling efficient collaboration between various grid components without compromising data confidentiality, FedGrid helps optimize energy distribution and usage. The approach also addresses the key challenges associated with data sharing and collaboration in smart grids, ultimately improving the overall efficiency and reliability of the system. As shown in Figure 3, energy optimization is crucial for coordinating multiple energy resources and consumption, especially when using FL techniques to protect data privacy.
Energy Management In (Abdellatif et al., 2024), a Secure, Distributed, and Collaborative Learning (SDCL) framework is presented to enhance the operations of smart grids. This framework focuses on maintaining data privacy and security while facilitating collaborative learning among grid components. By incorporating advanced AI and blockchain technologies, SDCL aims to optimize grid performance, improve system efficiency, and ensure reliable energy management. Additionally, (Lin and Ciou, 2024) introduces a privacy-preserving energy management framework that utilizes vertical FL for smart data cleaning in smart homes. This approach enhances user privacy while enabling effective analysis and management of smart meter data, ensuring that energy usage is optimized without compromising sensitive information.
@@ 3.2. Transmission and Distribution Stages
For energy distribution (Mourtzis et al., 2023), FL enables optimal energy distribution solutions and flexibility predictions, significantly improving the efficiency of power distribution in SGs. In the area of attack (Tran et al., 2023; Zhao et al., 2024) and anomaly detection (Kesici et al., 2024; Jithish et al., 2023), FL has demonstrated its effectiveness in identifying cyber attacks and anomalies within SG systems. This approach allows multiple stakeholders to collaboratively train robust detection models while maintaining data privacy, as raw consumption data remains local. In the realm of assessment (Li et al., 2024b; Ren et al., 2023b, a), FL has been applied to dynamic security assessment, enabling participants to collaboratively build models while preserving data privacy.
Energy Distribution Mourtzis et al. (Mourtzis et al., 2023) develop an integrated FL framework aims at enhancing the efficiency and security of energy distribution in smart grids. This framework addresses privacy concerns by enabling collaborative learning without sharing raw data, which is particularly relevant in smart grid data analysis. Specifically, they propose an incentive mechanism for FL, edge devices are compensated by service providers based on their contribution to a given task. The payoff function for an edge device $i$ participating in task $\Upsilon_{j,k}$ under service provider $j$ is:
@@
Attack Detection Tran et al. (Tran et al., 2023) introduce a privacy-preserving FL approach specifically designed to detect false data injection attacks (FDIAs) in smart grid systems. This method focuses on enhancing data privacy and security in cross-silo FL scenarios within the SG environment. By improving the detection efficiency and effectiveness of FDIAs, their approach ensures that data confidentiality is maintained across various entities involved in the grid, making it a robust solution for securing SG infrastructure. Zhao et al. (Zhao et al., 2024) propose a FL algorithm that integrates client sampling and gradient projection techniques to enhance SG applications. This approach effectively addresses privacy concerns by enabling model training using fragmented data from devices spread across the grid. Furthermore, the inclusion of client sampling and gradient projection techniques leads to improved FL performance and optimization, providing a more reliable framework for handling SG data without compromising security.
@@
Anomaly Detection Kesici et al. (Kesici et al., 2024) propose a data-driven time series anomaly detection method that leverages time-aware shapelets, a shapelet evolution graph, and segment embeddings learned through the DeepWalk algorithm. Their method demonstrates a significant improvement in detecting FDIA, with a 20-40% increase in accuracy compared to other unsupervised learning techniques. Additionally, when applied with BO-XGBoost, the model achieves a 5% higher accuracy than unmodified XGBoost, adding further robustness to the detection process. Moreover, this approach imparts physical significance to the dynamic evolution of time series models, providing valuable insights into the underlying data. Jithish et al. (Jithish et al., 2023) present a FL-based approach for distributed anomaly detection in SGs. Their solution enhances privacy and security by allowing for collaborative detection of anomalies without the need to centralize sensitive data. This distributed approach not only safeguards data privacy but also improves the overall stability and reliability of SG systems. By enabling effective anomaly detection across the grid, their method ensures that potential issues are identified early, contributing to more resilient SG operations.
@@
Security Assessment Li et al. (Li et al., 2024b) introduce a quantum-secured FL system, known as QFDSA, for dynamic security assessment in SGs. This system integrates quantum-resistant cryptography with FL to ensure both data privacy and security, particularly in the context of SG dynamic security assessments (DSA). Their approach addresses the limitations of traditional DSA methods, which may be vulnerable to privacy breaches and quantum computing attacks, thereby offering a more secure alternative for assessing grid security. Ren et al. (Ren et al., 2023b) further develop QFDSA, a hybrid system that combines quantum security and FL for dynamic security assessment in SGs. By enhancing privacy and security measures, their approach tackles key challenges related to data protection and collaborative analysis in SG environments. Specifically, in Quantum Key Distribution (QKD) for FL, secure keys are exchanged between participants, typically denoted as Alice and Bob, using quantum states to ensure security. In a Measurement-Device-Independent QKD (MDI-QKD) protocol, both Alice and Bob randomly choose a measurement basis from either $K$ or $Z$ and send weak coherent-state pulses (WCPs) with intensities sampled from predefined sets. These pulses are transmitted to a trusted or untrusted relay (Charlie) which performs a Bell state measurement (BSM). The key rate $R$ for secure communication is calculated by:
@@
Ren et al. (Ren et al., 2023a) propose EFedDSA, an efficient differential privacy (DP)-based horizontal FL approach for SG dynamic security assessment. This system enhances privacy protection in data-driven security assessments, addressing the challenge of data isolation and privacy concerns in traditional machine learning approaches. EFedDSA improves the efficiency and effectiveness of dynamic security assessment while maintaining data confidentiality across multiple participants in the SG infrastructure, offering a comprehensive solution for modern SG security needs.

### v2

@@ 1. Introduction
The Smart Grid (SG), which integrates the physical infrastructure of the electrical grid with modern information and communication technologies (ICTs) (vasisht2022smart), is a prime example of a cyber-physical system (CPS). Unlike conventional power grids, SGs can detect, respond to, and proactively manage changes in electricity demand and supply, while also addressing various operational challenges (obeagu2022overview).
The SG architecture incorporates several key components that work together to enhance the overall functionality of the electrical system (mittal2024smart), including advanced metering infrastructure (AMI) (molokomme2019survey), communication networks (hu2023multi), and sophisticated information technology solutions (reza2014overview), among others. By leveraging real-time data and advanced analytics, SGs aim to improve the efficiency of power generation, transmission and distribution, and consumption (chopda2021overview). Although SGs enable a more flexible and resilient power system that can better accommodate the growing demand for electricity, the increasing reliance on ICTs has made SG systems more vulnerable to security and privacy attacks. Federated Learning (FL) has emerged as a promising approach to addressing privacy and security concerns in SG systems. This innovative distributed machine learning paradigm enables collaborative model training across multiple participants without the need to share raw data, making it particularly suitable for SG applications, where generators are distributed and data privacy for each generator is crucial. We categorize the application of FL in SGs into three aspects: generation, transmission and distribution, and consumption. From the perspective of generation (tuballa2016review) in SGs, FL is primarily being used to enhance renewable energy integration and forecasting (camacho2011control). FL enables distributed SG generators to collaboratively train models for predicting resources such as solar and wind power output without sharing sensitive operational data. In the transmission (bayoumi2015power) and distribution aspects (hamidi2010smart) of SGs, FL is proving valuable for line fault detection and classification, optimal power flow (usman2022novel), voltage control (xie2021review), and more. By utilizing a FL-based approach, grid operators can quickly identify and locate faults in transmission lines, while collaboratively training models for power distribution without compromising the privacy of individual generators. In the consumption process of SGs (kaur2020energy), FL is extensively used for load forecasting (husnoo2023secure) and energy theft detection (li2024fuse). Smart meters can collaboratively train models to predict household energy consumption patterns without sharing individual usage data.
@@
Grataloup et al. (grataloup2024review) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zheng et al. (zheng2024advancing) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zhang et al. (zhang2024vulnerability) 2024 $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Pandya et al. (pandya2023federated) 2023 $\surd$ $\surd$ $\surd$ $\surd$
Cheng et al. (cheng2022review) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Mirzaee et al. (mirzaee2022smart) 2022 $\surd$ $\surd$ $\surd$
Ding et al. (ding2022cyber) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Cui et al. (cui2020detecting) 2020 $\surd$ $\surd$ $\surd$ $\surd$
@@ 1.1. Related Surveys
Grataloup et al. (grataloup2024review) focus on the integration of FL to address challenges in renewable energy applications, such as energy management, system reliability, energy consumption prediction, and load forecasting. However, this survey only highlights FL use cases in SGs as privacy-preserving collaborative learning approaches and overlooks the potential vulnerabilities of SGs when using deep learning and federated learning technologies. Zheng et al. (zheng2024advancing) provide a detailed discussion of components in FL systems, including model aggregation, data partitioning, and local model training. However, their survey paper lacks a discussion on model-related attacks within FL-based SG systems. Zhang et al. (zhang2024vulnerability) offer a comprehensive survey of applications for each stage of SG systems and discuss potential privacy and security violations. However, their focus is primarily on centralized machine learning algorithms. Our work bridges the gaps between SG, FL, and privacy and security vulnerabilities. Pandya et al. (pandya2023federated) demonstrate a broader scope, covering smart cities with a FL framework to preserve privacy. This approach enables effective utilization of large-scale data in areas such as electric supply chain monitoring, air quality forecasting, water consumption prediction, traffic management, and pandemic diagnosis. However, this work lacks details on SG applications and security issues within FL frameworks. Cheng et al. (cheng2022review) conduct a survey of the FL paradigm, comparing the types, network structures, and global model aggregation methods in SG. However, the brief survey lacks a discussion on robustness. Mirzaee et al. (mirzaee2022smart) and Ding et al. (ding2022cyber) discuss conventional security threats to SGs as well as emerging issues related to the application of machine learning techniques. They cover various types of cyberattacks, privacy concerns related to smart meter data, and potential countermeasures to address these challenges. However, they only cover a limited number of applications in SGs and primarily focus on centralized cases. Cui et al. (cui2020detecting) mainly focus on false data attacks and privacy issues in SGs. Compared to previous literature, our paper provides a comprehensive discussion on FL-based SGs and their potential vulnerabilities. We hope this can complement existing research and inspire further studies in this direction.
@@ 2.1. Smart Grid
The electric power grid has evolved by leaps and bounds since its inception (masera2018smart). A major development in this domain was the integration of ICT. The integration of ICT with the traditional power grid facilitated distributed power generation via better resource management and automated control (simoes2012comparison). ICT integration also marked the evolution of the conventional grid into the SG and solved several limitations including unidirectional power flow and centralized bulk generation of electricity. Additionally, it reduced the power and energy domain’s dependence on fossil fuels like coal and natural gas, leading the way for sustainability and decarbonization via the development of microgrids (jahromi2023cyber; yadav2020microgrid). The SG power flow framework consists of the following stages:
@@
    Generation: This stage refers to the production of AC/DC power. In the conventional power grid, power generation was centralized in nature. However, with ICT integration for real-time monitoring, power generation can also be achieved in a regulated manner at residential buildings (for example using solar photovoltaic panels) and used for local consumption/fed back to the grid (kahrobaee2012multiagent).
@@
    Transmission/Distribution: This stage refers to the transfer of the power produced at the generation stage to the appropriate end nodes for utilization. This involves the use of transformers, transmission lines, substation equipment and other line devices like fault detector relays, breakers to ensure a safe and reliable transfer at appropriate bus voltage levels. The fault detector relays and breakers are often digital in nature and controlled by the central monitoring station via ICT devices (gaspar2023smart).
@@
    Consumption: This is final stage of the SG power flow process and involves the utilization of the power as received from the transmission and distribution networks for powering electrical devices. The consumption level can involve entities like government agencies, private industries, private individuals, etc., each of which can have a different consumption footprint. In the modern SG, integration of ICT devices means that some consumer-level devices like electric vehicles can also be used to transfer surplus/unused power back to the grid (coordinated via V2G communications (madawala2011bidirectional)).
@@
While ICT integration in the power grid enabled easier and more robust resource management from a single centralized Supervisory Control and Data Acquisition (SCADA)-based control system (gaushell1993scada), the increasing number of IoT devices and local generation facilities like solar panels means that it can be extremely computationally resource intensive to monitor them. Inadequate monitoring facilities can mean that intrusive events like cyberattacks and line faults may not be detected and mitigated on time.
@@
Slowed mitigation of cyberattacks, faults and other control emergencies can have a devastating impact on the stability of the SG (jahromi2023cyber). Moreover, the SG being a cyber-physical system implies that cyber layer vulnerabilities can be exploited to cause severe physical ramifications. For example, cyberattacks on the SG can lead to uncontrolled voltage increase at one or more buses leading to equipment damage and consumer-level device burnout (rath2022behind). To alleviate these issues, several researchers have proposed, implemented and evaluated the use of Artificial Intelligence (AI)-based tools and techniques for applications like real-time control (bose2017artificial), cybersecurity (ruan2023super), load forecasting (zhou2017holographic), and fault detection (sun2023high). These AI tools are typically deployed at the centralized controller level for monitoring, removing the need for human system operators.
@@
As the AI modules employed for applications like cybersecurity/control monitoring and load forecasting in the SG environment, are often placed at a central server, they require the exchange of data between local Distributed Energy Resources (DERs) and the central node via communication networks. However, this presents privacy risks (sun2021solar) as the data being transferred can be utilized to reveal energy consumption patterns of private individuals (liu2018practical; zhao2020smart; wang2020spds). This is a major risk as the consumption patterns can be cross-analyzed to determine when the individuals are at home/outside. Further, their home appliances can also be revealed motivating malicious entities to schedule burglary attempts (liu2018practical).
@@
Another risk that the communication-based training data exchange between local DERs and the centralized server housing AI modules is the non-willingness within third-party DER owners to share proprietary data. This non-willingness typically stems from the concern that the proprietary data may be accessed by malicious entities who can then know the system details of the DER. This can be used to manipulate the system via stealth cyberattack vectors (leng2021stability). Non-willingness can also arise from a lack of trust between the DER owners and the utility company that manages the SG and the control center.
@@
Both the above-mentioned factors can lead to data constraints that may severely hinder the AI training process leading to flawed and untrustworthy decision-making, raising concerns on the reliability of AI. To solve the problem, An AI training strategy called Federated Learning can be used. In this strategy proprietary training data does not need to be transferred over a communication channel leading to alleviation of privacy concerns (mcmahan2017communication). More details about this training strategy are available in the following subsection.
@@ 2.2. Federated Learning
To solve the above optimization problem, the system runs for $T$ rounds of FL training. Initially, the server stores the global model $\mathbf{w}^{0}$. In round $t$, the server randomly selects a subset $\mathcal{Q}^{t}$ of $q$ devices and sends them the latest global model $\mathbf{w}^{t}$. These selected devices update the global model $\mathbf{w}^{t}$ using their local data. For instance, in the widely-used Federated Averaging (FedAvg) (fl_original) algorithm, each selected device performs $\kappa$ iterations of stochastic gradient descent (SGD) to update the local model as follows:
@@ 3.1. Generation Stage
FL enables collaborative model training across different energy producers and consumers, enhancing the accuracy of energy generation forecasts and enabling dynamic resource allocation without the need to centralize sensitive operational data (lee2023towards; gupta2023fedgrid). In the context of energy management (abdellatif2024sdcl), FL supports the development of intelligent systems that can predict and optimize power generation patterns while keeping sensitive user data local on smart devices. This approach improves grid efficiency and facilitates more effective demand response programs.
@@
Energy Forecasting In the work by (lee2023towards), an advanced algorithm for FL is proposed to tackle the challenges of non-IID data in distributed solar energy grids. This framework is specifically designed to handle the heterogeneous nature of the data across different solar energy producers, enhancing the model’s ability to converge efficiently in decentralized settings. By addressing the variations in data distribution, the framework improves the accuracy and reliability of predictions, leading to better performance in solar energy generation and more effective resource allocation.
Energy Optimization The FedGrid framework introduced by (gupta2023fedgrid) offers a secure FL solution for optimizing energy management within smart grids. This framework emphasizes the importance of protecting privacy and enhancing data security during the optimization process. By enabling efficient collaboration between various grid components without compromising data confidentiality, FedGrid helps optimize energy distribution and usage. The approach also addresses the key challenges associated with data sharing and collaboration in smart grids, ultimately improving the overall efficiency and reliability of the system. As shown in Figure 3, energy optimization is crucial for coordinating multiple energy resources and consumption, especially when using FL techniques to protect data privacy.
Energy Management In (abdellatif2024sdcl), a Secure, Distributed, and Collaborative Learning (SDCL) framework is presented to enhance the operations of smart grids. This framework focuses on maintaining data privacy and security while facilitating collaborative learning among grid components. By incorporating advanced AI and blockchain technologies, SDCL aims to optimize grid performance, improve system efficiency, and ensure reliable energy management. Additionally, (lin2024privacy) introduces a privacy-preserving energy management framework that utilizes vertical FL for smart data cleaning in smart homes. This approach enhances user privacy while enabling effective analysis and management of smart meter data, ensuring that energy usage is optimized without compromising sensitive information.
@@ 3.2. Transmission and Distribution Stages
For energy distribution (mourtzis2023integration), FL enables optimal energy distribution solutions and flexibility predictions, significantly improving the efficiency of power distribution in SGs. In the area of attack (tran2023efficient; zhao2024federated) and anomaly detection (kesici2024detection; jithish2023distributed), FL has demonstrated its effectiveness in identifying cyber attacks and anomalies within SG systems. This approach allows multiple stakeholders to collaboratively train robust detection models while maintaining data privacy, as raw consumption data remains local. In the realm of assessment (li2024privacy; ren2023qfdsa; ren2023efeddsa), FL has been applied to dynamic security assessment, enabling participants to collaboratively build models while preserving data privacy.
Energy Distribution Mourtzis et al. (mourtzis2023integration) develop an integrated FL framework aims at enhancing the efficiency and security of energy distribution in smart grids. This framework addresses privacy concerns by enabling collaborative learning without sharing raw data, which is particularly relevant in smart grid data analysis. Specifically, they propose an incentive mechanism for FL, edge devices are compensated by service providers based on their contribution to a given task. The payoff function for an edge device $i$ participating in task $\Upsilon_{j,k}$ under service provider $j$ is:
@@
Attack Detection Tran et al. (tran2023efficient) introduce a privacy-preserving FL approach specifically designed to detect false data injection attacks (FDIAs) in smart grid systems. This method focuses on enhancing data privacy and security in cross-silo FL scenarios within the SG environment. By improving the detection efficiency and effectiveness of FDIAs, their approach ensures that data confidentiality is maintained across various entities involved in the grid, making it a robust solution for securing SG infrastructure. Zhao et al. (zhao2024federated) propose a FL algorithm that integrates client sampling and gradient projection techniques to enhance SG applications. This approach effectively addresses privacy concerns by enabling model training using fragmented data from devices spread across the grid. Furthermore, the inclusion of client sampling and gradient projection techniques leads to improved FL performance and optimization, providing a more reliable framework for handling SG data without compromising security.
@@
Anomaly Detection Kesici et al. (kesici2024detection) propose a data-driven time series anomaly detection method that leverages time-aware shapelets, a shapelet evolution graph, and segment embeddings learned through the DeepWalk algorithm. Their method demonstrates a significant improvement in detecting FDIA, with a 20-40% increase in accuracy compared to other unsupervised learning techniques. Additionally, when applied with BO-XGBoost, the model achieves a 5% higher accuracy than unmodified XGBoost, adding further robustness to the detection process. Moreover, this approach imparts physical significance to the dynamic evolution of time series models, providing valuable insights into the underlying data. Jithish et al. (jithish2023distributed) present a FL-based approach for distributed anomaly detection in SGs. Their solution enhances privacy and security by allowing for collaborative detection of anomalies without the need to centralize sensitive data. This distributed approach not only safeguards data privacy but also improves the overall stability and reliability of SG systems. By enabling effective anomaly detection across the grid, their method ensures that potential issues are identified early, contributing to more resilient SG operations.
@@
Security Assessment Li et al. (li2024privacy) introduce a quantum-secured FL system, known as QFDSA, for dynamic security assessment in SGs. This system integrates quantum-resistant cryptography with FL to ensure both data privacy and security, particularly in the context of SG dynamic security assessments (DSA). Their approach addresses the limitations of traditional DSA methods, which may be vulnerable to privacy breaches and quantum computing attacks, thereby offering a more secure alternative for assessing grid security. Ren et al. (ren2023qfdsa) further develop QFDSA, a hybrid system that combines quantum security and FL for dynamic security assessment in SGs. By enhancing privacy and security measures, their approach tackles key challenges related to data protection and collaborative analysis in SG environments. Specifically, in Quantum Key Distribution (QKD) for FL, secure keys are exchanged between participants, typically denoted as Alice and Bob, using quantum states to ensure security. In a Measurement-Device-Independent QKD (MDI-QKD) protocol, both Alice and Bob randomly choose a measurement basis from either $K$ or $Z$ and send weak coherent-state pulses (WCPs) with intensities sampled from predefined sets. These pulses are transmitted to a trusted or untrusted relay (Charlie) which performs a Bell state measurement (BSM). The key rate $R$ for secure communication is calculated by:
@@
Ren et al. (ren2023efeddsa) propose EFedDSA, an efficient differential privacy (DP)-based horizontal FL approach for SG dynamic security assessment. This system enhances privacy protection in data-driven security assessments, addressing the challenge of data isolation and privacy concerns in traditional machine learning approaches. EFedDSA improves the efficiency and effectiveness of dynamic security assessment while maintaining data confidentiality across multiple participants in the SG infrastructure, offering a comprehensive solution for modern SG security needs.